OpenStudy (anonymous):

Given f(x) = 5x - 6 and g(x) = x2 - 5x + 6, find the function (f - g)(x)

OpenStudy (anonymous):

\[(f-g)(x)=f(x)-g(x)=5x-6-(x^2-5x+6)=-x^2+10x-12\]

OpenStudy (anonymous):

its not good to just give them the answer you know... they need to try them selves

OpenStudy (anonymous):

subtracting f(x)-(g) 6 thats that and idk how to do it?...

OpenStudy (anonymous):

thanks thomas9 helped alot

OpenStudy (anonymous):

So do you understand it?

OpenStudy (anonymous):

yes but i dk if i got it right though ha